OpenStudy (ninaa.):

What is the simplest form of 82 over 100?

OpenStudy (abb0t):

\(\sf \frac{82}{100} = \huge \frac{\frac{82}{2}}{\frac{100}{2}}\), now if you divide each by 2, what do you get?

OpenStudy (anonymous):

\[82 \over 100\]

OpenStudy (anonymous):

or 41/50

OpenStudy (abb0t):

Maybe you can explain to them how you got that, @RatzLover95

OpenStudy (the_fizicx99):

\[\frac{ 82 }{ 100 }\] dividing both by 2 will give you \[82 \div 2 = 41\] \[100 \div 2 = 50\] \[\frac{ 41 }{ 50 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

right ^^^^

OpenStudy (anonymous):

ya